OpenStudy (anonymous):

Fill in the blank with the numerator of the difference shown below. x+2/x-3 - x+1/x-3

OpenStudy (anonymous):

\(\huge \color{blue}{\textbf{Welcome To Openstudy..}}\)

OpenStudy (anonymous):

=x+2/x-3 - x+1/x-3 =x+2 -(x+1)/x-3 =x+2-x-1/x-3 =2-1/x-3 =1/x-3

OpenStudy (anonymous):

thank you so much :))

OpenStudy (anonymous):

You must give the medal to Qibtiya and not me... I have just welcome you..

OpenStudy (anonymous):

@waterineyes =)

OpenStudy (anonymous):

@waterineyes @danay thank u =)

OpenStudy (anonymous):

Oh I forgot that even I can give the medal.. Sorry..

OpenStudy (anonymous):

@waterineyes its ok =)...and thank u once again

OpenStudy (anonymous):

\(\huge \color{green}{\textbf{Welcome Dear..}}\)

OpenStudy (anonymous):

@waterineyes =)